Question 1
The profit made by a company when 60 units of its product is sold is R 1 600.00. When 150 units of its products are sold, the profit increases to R 5 200.00. Assuming that the profit function is linear find its equation.
1
Expert's answer
2019-05-02T10:04:31-0400

Let the desired linear function be

"y=kx+b"

where x is number of units, y is the profit. We need to find k and b. Make a system of equations.


"1600=60k+b""and""5200=150k+b"

Then

"3600=90k""k=40"

Then


"b=5200-40*150=-800"

Then

"y=40x-800"

Answer: y=40x-800